OpenStudy (anonymous):

Rachel purchased different carpets that had the same ratio of the length to the width. The table below shows the length and width of some carpets Rachel bought: Length (feet) Width (feet) 4 3 8 6 12 9 16 12 ? 15

OpenStudy (anonymous):

@Rushwr

OpenStudy (anonymous):

@Missiey

OpenStudy (anonymous):

The missing value in the table is ______ feet.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (rushwr):

Wait I think the table of lengths and widths are a little unclear . Can u show it in separate columns.

OpenStudy (anonymous):

4,8,12,16,? is length feet

OpenStudy (anonymous):

3,6,9,12,15 is width feet

OpenStudy (anonymous):

@Rushwr

OpenStudy (anonymous):

@pooja195

OpenStudy (rushwr):

If u go thru the 2 table thoroughly u can see a pattern in both cases. Can u identify them ?

OpenStudy (anonymous):

no

OpenStudy (rushwr):

okai look at the table that has the no. for length. 4,8,16,12, ? 4*1 =4 4*2 =8 4*3 =12 Right?

OpenStudy (rushwr):

Now look at the width table 3,6,9,12,15 3*1 =3 3*2=6 3*3=9 3*4=12 3*5=15

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so is it 20

OpenStudy (rushwr):

yes !!!!! Good job !

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i have a few more will u help

OpenStudy (rushwr):

Ur welcome :)

OpenStudy (rushwr):

Okai ! I'll be a little late to respond though.

OpenStudy (anonymous):

ok